OpenStudy (anonymous):

Please help! Use the formula s=y theta to solve the following: The radius of a wheel is 10 inches. How far does the wheel travel as it turns through one-fourth revolution? a. 6.4 in b. 15.7 c. 12.7 d. 7.9

OpenStudy (anonymous):

r = 10 inches.. \(Circumference = 2 \pi r\) So, In one revolution C = 2*(3.14)*(10) = 62.8 Now, in one fourth revolution, \(\frac{1}{4} \times 62.8 = 15.7 \)

OpenStudy (anonymous):

Or this can be done in one more way:

OpenStudy (anonymous):

\(s= y. \theta\) \(\theta = \frac{\pi}{2}\) here.. \(y = 10\) So, \[s = y.\theta\] \[s= 10 \times \frac{\pi}{2} \implies 10 \times 1.57 \implies 15.7\]

OpenStudy (anonymous):

Oh I see. Thank you so much. I really appreciate the help

OpenStudy (anonymous):

Welcome dear..